3.3 g of magnesium oxide are produced when 3.0 g of magnesium reacts with 5.0 g of oxygen gas. Calculate the percent yield of magnesium oxide produced.


1
Expert's answer
2022-04-28T02:40:03-0400

2Mg + O2 = 2MgO

n(Mg) = 3.0/24 = 0.125 mol -- limiting reactant

n(O2) = 5.0/32 = 0.15625 mol

n(MgO)theor = n(Mg) = 0.125 mol

n(MgO)real = 3.3/(24+16) = 0.0825 mol

yield = 0.0825/0.125 = 0.66 = 66%